OpenStudy (anonymous):

**WILL MEDAL** Kara graphed the system of equations and stated that the solution is (4, -2). Check her solution algebraically. Which statement is true? A. The graph of y=-x+2 is incorrect. B. The graph of y=1/2x-4 is incorrect. C. (-2, 4) is the solution to the system of equations. D. (4,-2) is the solution to the system of equations.

OpenStudy (anonymous):

A) -2 = -(4) +2 -2= =2 A B) -2 = 1/2(4) -4